OpenStudy (anonymous):

12. What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form. (1 point) y = 4x y = –4x y = –1x – 4 y = –4x + 1

OpenStudy (anonymous):

-1-1/-4-4= 2/8=1/4 Y=mx+b y= 1/4x+4

OpenStudy (anonymous):

wait so whats the answer? Thats not a choice?

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

y = 4x y = –4x y = –1x – 4 y = –4x + 1

OpenStudy (anonymous):

oh duh..i see them now hold on

OpenStudy (anonymous):

Y= 4x Sorry I wrote it wrong last time

OpenStudy (anonymous):

thanks you so much! :DD

OpenStudy (anonymous):

you're welcome